# Hermite trace polynomials and chaos decompositions for the Hermitian Brownian motion

### v2

@@ Abstract.
For a non-zero parameter $q$, we define Hermite trace polynomials, which are multivariate polynomials indexed by permutations. We prove several combinatorial properties for them, such as expansions and product formulas. The linear functional determined by these trace polynomials is a state for $q=\frac{1}{N}$ for $N$ a non-zero integer. For such $q$, Hermite trace polynomials of different degrees are orthogonal. The product formulas extend to the closure with respect to the state. The state can be identified with the expectation induced by the $N\times N$ Hermitian Brownian motion. Hermite trace polynomials are martingales for this Brownian motion, while the elements in the closure can be interpreted as stochastic integrals with respect to it. Using the grading on the algebra, we prove several chaos decompositions for such integrals, as well as analyse corresponding creation and annihilation operators. In the univariate, pure trace polynomial case, trace Hermite polynomials can be identified with the Hermite polynomials of matrix argument.
@@ 1. Introduction
  • Expansion: $H_{n}(x,t)=\sum_{j=0}^{[n/2]}(-1)^{j}\frac{n!}{(n-2j)!2^{j}j!}t^{j}x^{n-2j}$, where the coefficients are related to the number of incomplete matchings;
  • Product formulas: $H_{n}(x,t)H_{k}(x,t)=\sum_{j=0}^{n\wedge k}\binom{n}{j}\binom{k}{j}j!t^{j}H_{n+k-2j}(x,t)$, where the coefficients are related to the number of inhomogeneous incomplete matchings;
@@
[…]
$\mathbb{E}\left[X(t)^{3}|\leq s\right]=X(s)^{3}+(t-s)\left(2X(s)+\frac{1}{N}\operatorname{\mathrm{Tr}}[X(s)]I_{N}\right).$
It is therefore natural to replace the algebra of polynomials by a larger algebra of trace polynomials. Here a trace monomial is, roughly speaking, a product of a regular monomial and traces of regular monomials. See Section 6.1 for details. An incomplete but representative list of related work involving trace polynomials includes the study of random matrices [Céb13], [DHK13, Kem16, Kem17], noncommutative functions [KŠ17, KPV21], and operator algebras [DGS21, Jek20, JLS22]. Other related work includes [GV06, Lév08, Hub21].
In this article, we study the $\ast$-algebra of trace polynomials, the state on it induced by the Hermitian Brownian motion, the corresponding basis of Hermite trace polynomials, and the larger algebra obtained as its completion. However, most of the results in the article are stated in terms of permutations (rather than trace monomials), a general parameter $q$ (rather than $\frac{1}{N}$), and a general Hilbert space $\mathcal{H}$ (rather than $L^{2}(\mathbb{R}_{+},dx)$). Thus, denote by $S_{0}(n)$ the symmetric group on the set $\left\{0,1,\ldots,n\right\}$. For each $n$, $\alpha\in S_{0}(n)$ and $h_{1},\ldots,h_{n}\in\mathcal{H}_{\mathbb{R}}$, we will consider symbols $\mathnormal{T}\left(\alpha\otimes(h_{1}\otimes\ldots\otimes h_{n})\right)$ subject to the symmetry relation
@@
Here for $\alpha\in S_{0}(n)$, $\beta\in S_{0}(k)$, the permutation $\alpha\cup\beta\in S_{0}(n+k)$ is obtained by shifting the cycles of $\beta$ by $n$, and merging the cycles of $\alpha$ and $\beta$ containing 00. See Notation 3.2.
Next, let $q$ be a non-zero parameter. For a transposition $\tau$, we define the contraction $C_{\tau}$ as follows. For $\alpha\in S_{0}(n)$, $C_{\tau}(\alpha)$ is obtained by: multiplying $\tau\alpha$; erasing the support of $\tau$ and shifting to obtain a permutation in $S_{0}(n-2)$; and multiplying by a weight depending on $q$ and the number of cycles in $\tau\alpha$. See Definition 3.9. Then as usual, we define the Laplacian as the sum over transpositions $\mathcal{L}=\sum_{\tau}C_{\tau}$, and the Hermite trace polynomial $\mathnormal{W}\left(\eta\otimes_{s}F\right)=\mathnormal{T}\left(e^{-\mathcal{L}}(\eta\otimes_{s}F)\right)$. Here the contraction on the tensor part of the argument is the usual tensor contraction. Hermite trace polynomials satisfy several properties which parallel those of ordinary Hermite polynomials.
We may now define a linear functional $\varphi$ on $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ by requiring it to be unital and zero on each $\mathnormal{W}\left(\eta\otimes_{s}F\right)$. If we use $\left\{\mathnormal{T}\left(\eta\otimes_{s}F\right)\right\}$ as a spanning set for $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$, multiplication does not depend on $q$, while the state $\varphi$ does. On the other hand, if we use $\left\{\mathnormal{W}\left(\eta\otimes_{s}F\right)\right\}$ as a spanning set, the state does not depend on $q$, but multiplication does. $\varphi$ is positive semi-definite if and only if $q$ is zero or of the form ±$\pm\frac{1}{N}$ for $N\in\mathbb{N}$. For such $q$, we can define the GNS Hilbert space $\mathcal{F}_{q}(\mathcal{H})$ as the completion of the quotient $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$ of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ with respect to the seminorm induced by $\varphi$. The inner product takes the form
@@
Since $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ is naturally graded, we may interpret $\mathcal{F}_{q}(\mathcal{H})$ as a Fock space. There are several natural choices of creation (and annihilation) operators which we describe. The corresponding field operators live in smaller subalgebras of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$. There is an interesting connection with a different Fock space construction from [BG02] which we describe in some detail. Incidentally, another article exploring the connection between characters of the symmetric groups and GUE matrices is [KN21]. We have not elucidated the connection between their work and ours.
@@
We thus may interpret Hermite trace polynomials as elements $\mathnormal{W}\left(\alpha\otimes_{s}(h_{1}\otimes\ldots\otimes h_{n})\right)$ in the algebra $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$ with the state $\varphi$; or the corresponding trace polynomials in formal variables; or (for $q=1/N$) the corresponding trace polynomials in Gaussian random matrices $\left\{X(h_{i})\right\}$. In this last interpretation, we prove that the matrix entries of Hermite trace polynomials are themselves (multivariate) Hermite polynomials in the matrix entries of $\left\{X(h_{i})\right\}$.
To the best of our knowledge, the Hermite trace polynomials $\mathnormal{W}\left(\eta\otimes_{s}F\right)$ are new even in the univariate case $\mathcal{H}=\mathbb{C}$. However if we further restrict to $\eta\in\mathbb{C}[S(n)]$ (rather than $S_{0}(n)$), the corresponding objects have appeared in the literature. Indeed, denoting $\chi^{\lambda}$ the character of the irreducible representation indexed by the partition $\lambda$, $\mathnormal{W}\left(\chi^{\lambda}\right)$ is closely related to the corresponding Hermite polynomial of matrix argument. In particular, these elements are all orthogonal with respect to $\varphi$. Moreover, one can form a more general set of characters $\chi^{\lambda,\lambda^{\prime}}$ indexed by pairs of partitions which differ by one box, such that $\left\{\mathnormal{W}\left(\chi^{\lambda,\lambda^{\prime}}\right)\right\}$ form an orthogonal basis for $\mathcal{F}_{1/N}(\mathbb{C})$. This collection of trace polynomials is clearly deserving of additional study.
@@
Acknowledgements. The first author has discussed various aspects of this project with a number of people. He is especially grateful to Todd Kemp, Andu Nica, JM Landsberg, Marek Bożejko, and Jurij Volčič. The authors are also grateful to the referee(s), whose comments resulted in substantial improvements, particularly the Gaussian Hilbert spaces approach and Theorem 6.11.
@@ 2.1. Permutations and partitions
Denote $[n]=\left\{1,\ldots,n\right\}$ and $[0,n]=\left\{0,1,\ldots,n\right\}$.
@@
Denote by $S_{0}(n)$ the permutations of $[0,n]$. $S(n)$ is a subgroup of $S_{0}(n)$, which acts on it by conjugation. The equivalence classes under this action are subsets of the standard conjugacy classes where the number of elements of the cycle containing 00 is preserved. So they are in a natural bijection with number partitions of $n+1$ with a marked element. Equivalently, they are indexed by pairs of partitions $(\lambda,\lambda^{\prime})$, where $\lambda\in\operatorname{\mathrm{Par}}(n)$, $\lambda^{\prime}\in\operatorname{\mathrm{Par}}(n+1)$, and the corresponding diagrams differ by one box.
For $m>n$, we will identify the element $\alpha\in S_{0}(n)$ with the corresponding element of $S_{0}(m)$ under the natural inclusion $[0,n]\subset[0,m]$.
For $\alpha\in S_{0}(n)$, denote $\operatorname{\mathrm{cyc}}_{0}(\alpha)=\operatorname{\mathrm{cyc}}(\alpha)-1$, where $\operatorname{\mathrm{cyc}}(\alpha)$ is the number of cycles of $\alpha$. In other words, $\operatorname{\mathrm{cyc}}_{0}(\alpha)$ is the number of cycles of $\alpha$ not containing 00. Denote
@@
To make the paper more readable, we will write the elements of the group algebra $\mathbb{C}[S_{0}(n)]$ as linear combinations $\sum_{\alpha\in S_{0}(n)}c_{\alpha}\alpha$ rather than the more standard $\sum_{\alpha\in S_{0}(n)}c_{\alpha}\delta_{\alpha}$.
A partial permutation $\alpha\in\mathcal{PS}_{0}(n)$ is a bijection from a subset of $[0,n]$ onto a subset of $[0,n]$; proper subsets and the empty subset are allowed. Orbits of a partial permutation fall into two types. Cyclic orbits are cycles in the usual permutation sense. Linear orbits have the initial and the final element. Note that a linear orbit has at least two elements. It is convenient to abuse the terminology and consider elements of $[0,n]$ which do not belong to any orbit of $\alpha$ as single-element linear orbits of $\alpha$. Denote $\mathcal{PS}_{0}(n,N)$ the set of partial permutations of $[0,n]$ with $N$ linear orbits.
@@ 2.2. Structure theory of the symmetric group
For a partition $\lambda\in\operatorname{\mathrm{Par}}(n+1)$, denote by $\chi^{\lambda}$ the character of the irreducible representation of $S_{0}(n)$ corresponding to $\lambda$. We will identify $\chi^{\lambda}$ with the element
@@
$Z(\mathbb{C}[S_{0}(n)]:\mathbb{C}[S(n)])=\left\{\eta\in\mathbb{C}[S_{0}(n)]:{\sigma^{-1}}\eta\sigma=\eta\text{ for all }\sigma\in S(n)\right\}.$
The following are well-known [VOk04], [Gil05].
@@
  • For $\lambda\in\operatorname{\mathrm{Par}}(n)$, write $\lambda^{\prime}=\lambda+\square$ if the Young diagram for $\lambda^{\prime}$ is obtained by adding one box to the Young diagram for $\lambda$. Denote $\chi^{\lambda^{\prime}:\lambda}$ the character of the compression of the $\lambda^{\prime}$-irreducible representation of $S_{0}(n)$ to the (unique) component giving a $\lambda$-irreducible representation of $S(n)$. Then $\left\{\chi^{\lambda^{\prime}:\lambda}:\lambda\in\operatorname{\mathrm{Par}}(n),\lambda^{\prime}=\lambda+\square\right\}$ are orthogonal and span $Z(\mathbb{C}[S_{0}(n)]:\mathbb{C}[S(n)])$.
@@
See for example [BMP16] for a detailed exposition and explicit expressions for $\mathcal{W}(E_{ij}^{\lambda})$.
@@ 2.4. Gaussian Hilbert spaces
See [Jan97] for proofs and more details.
@@
and $\mathcal{P}(\mathcal{H}_{\mathbb{R}})=\bigcup_{n=0}^{\infty}\mathcal{P}_{n}(\mathcal{H}_{\mathbb{R}})=$ all such polynomials. The complexification $\mathcal{H}$ of $\mathcal{H}_{\mathbb{R}}$ is then naturally isomorphic to $L^{2}(\mathcal{P}(\mathcal{H}_{\mathbb{R}}),\mathbb{E})$. Denote $\overline{\mathcal{P}}_{n}(\mathcal{H}_{\mathbb{R}})$ the closure of $\mathcal{P}_{n}(\mathcal{H}_{\mathbb{R}})$ in $L^{2}(\mathcal{P}(\mathcal{H}_{\mathbb{R}}),\mathbb{E})$ and $\mathcal{H}_{n}=\overline{\mathcal{P}}_{n}(\mathcal{H}_{\mathbb{R}})\ominus\overline{\mathcal{P}}_{n-1}(\mathcal{H}_{\mathbb{R}})$, the orthogonal complement taken in $L^{2}(\mathcal{P}(\mathcal{H}_{\mathbb{R}}),\mathbb{E})$. Denote $p\mapsto:\!\!p\!\!:$ the orthogonal projection from each $\overline{\mathcal{P}}_{n}(\mathcal{H}_{\mathbb{R}})$ onto $\mathcal{H}_{n}$; note that it maps each $\mathcal{P}_{n}(\mathcal{H}_{\mathbb{R}})$ to itself. Define the Wick product map $W$ from the algebraic symmetric Fock space $\bigoplus_{n=0}^{\infty}\mathcal{H}_{\mathbb{R}}^{\odot_{s}n}$ into $L^{2}(\mathcal{P}(\mathcal{H}_{\mathbb{R}}),\mathbb{E})$ by the linear extension of
@@
$\mathnormal{W}\left(f_{1}\otimes\ldots\otimes f_{n}\right)=\sum_{\pi\in\mathcal{P}_{1,2}(n)}(-1)^{n-\left|\pi\right|}\prod_{\left\{i,j\right\}\in\mathrm{Pair}\left(\pi\right)}\left\langle f_{i},f_{j}\right\rangle\prod_{k\in\mathrm{Sing}\left(\pi\right)}X(f_{k}).$
@@
In the complex case?
There are two natural types of complex Gaussian Hilbert space.
  • We may take =$\mathcal{H}=\left\{f+ig:f,g\in\mathcal{H}_{\mathbb{R}}\right\}$ to be the complexification of $\mathcal{H}_{\mathbb{R}}$, with $X(f+ig)=X(f)+iX(g)$. Then each $X(h)$ is centered complex Gaussian, and their covariance is $\mathbb{E}[X(h)X(k)]=\left\langle h,\bar{k}\right\rangle_{\mathcal{H}}$.
@@
If $\left\{\xi_{i}:i\in\Xi\right\}$ is an orthonormal basis for $\mathcal{H}_{\mathbb{R}}$, then $\left\{\mathnormal{W}\left(\xi_{u(1)},\ldots,\xi_{u(n)}\right):n\geq 0,\mathbf{u}\in\Xi^{n}\right\}$ is an orthogonal basis for $L^{2}(\mathcal{P}(\mathcal{H}_{\mathbb{R}}),\mathbb{E})$. Similarly, $\left\{Z(\xi_{u(1)})\ldots Z(\xi_{u(n)}):n\geq 0,\mathbf{u}\in\Xi^{n}\right\}$ is an orthogonal basis for $L^{2}(\mathcal{P}(\mathcal{K}),\mathbb{E})$. Thus the linear extension of the map $\mathcal{S}:\mathnormal{W}\left(h_{1},\ldots,h_{n}\right)\mapsto Z(h_{1})\ldots Z(h_{n})$ (the Segal-Bargmann transform) is an isomorphism between these spaces.

### v3

@@ Abstract.
For a non-zero parameter $q$, we define Hermite trace polynomials, which are multivariate polynomials indexed by permutations. We prove several combinatorial properties for them, such as expansions and product formulas. The linear functional determined by these trace polynomials is a state for $q=\frac{1}{N}$ for $N$ a non-zero integer. For such $q$, Hermite trace polynomials of different degrees are orthogonal. The product formulas extend to the closure with respect to the state. The state can be identified with the expectation induced by the $N\times N$ Hermitian Brownian motion. Hermite trace polynomials are martingales for this Brownian motion, while the elements in the closure can be interpreted as stochastic integrals with respect to it. Using the grading on the algebra, we prove several chaos decompositions for such integrals, as well as analyze the corresponding creation and annihilation operators. In the univariate, pure trace polynomial case, trace Hermite polynomials can be identified with the Hermite polynomials of matrix argument.
@@ 1. Introduction
  • Expansion: $H_{n}(x,t)=\sum_{j=0}^{[n/2]}(-1)^{j}\frac{n!}{(n-2j)!2^{j}j!}t^{j}x^{n-2j}$, where the coefficients are related to the number of matchings of $n$ objects;
  • Product formulas: $H_{n}(x,t)H_{k}(x,t)=\sum_{j=0}^{n\wedge k}\binom{n}{j}\binom{k}{j}j!t^{j}H_{n+k-2j}(x,t)$, where the coefficients are related to the number of inhomogeneous matchings of $n$ objects;
@@
  • Stochastic integral representation: $H_{n}(B(t),t)=\int_{[0,t)^{n}}\,dB(t_{1})\ldots\,dB(t_{n})$.
Moreover, some of these results, notably the product formulas, extend to more general stochastic integrals
$\int_{[0,t)^{n}}f(t_{1},\ldots,t_{n})\,dB(t_{1})\ldots\,dB(t_{n})$
for $f\in L^{2}$.
$\mathbb{E}\left[X(t)^{3}|\leq s\right]=X(s)^{3}+(t-s)\left(2X(s)+\frac{1}{N}\operatorname{\mathrm{Tr}}[X(s)]I_{N}\right),$
where the conditional expectation is taken onto the subalgebra generated by the matrix entries up to time $s$. Thus to have a family closed under conditional expectations, the algebra of polynomials needs to be replaced by a larger algebra of trace polynomials. Here a trace polynomial in some collection of variables $\left\{x_{i}:i\in S\right\}$ is, roughly speaking, a polynomial in these variables as well as in the values of a tracial functional $\operatorname{\mathrm{tr}}$ applied to these variables. See [8] for a formal definition using a universal property. For example, $x^{3}+2x+\operatorname{\mathrm{tr}}[x]$ (as above) and
(1) $x_{2}x_{4}\operatorname{\mathrm{tr}}[x_{1}x_{3}x_{7}]\operatorname{\mathrm{tr}}[x_{5}x_{6}]$
are trace polynomials. See Section 6.1 for details. An incomplete but representative list of related work involving trace polynomials includes the study of random matrices [8, 12, 23, 24], noncommutative functions [27, 26], and operator algebras [10, 21, 22]. Other related work includes [35, 17, 30, 19].
In this article, we study the star-algebra of trace polynomials, the state on it induced by the Hermitian Brownian motion, the corresponding basis of Hermite trace polynomials, and the larger algebra obtained as its completion. However, we keep the context somewhat more general. First, a homogeneous trace monomial is naturally encoded by a permutation. Let $\alpha\in S_{0}(n)$, that is, a permutation of the set $\left\{0,1,\ldots,n\right\}$. Denote
@@
Compare with Notation 22.29 in [32]. For example, the trace monomial (1) corresponds to $\alpha=(024)(137)(56)$.
Second, the reciprocal matrix dimension $1/N$ is replaced by a general real parameter $q$. Finally, rather than considering variables indexed by time (or, extending linearly, by $L^{2}(\mathbb{R}_{+},dx)$), we index them by a general real Hilbert space $\mathcal{H}_{\mathbb{R}}$. Thus, for each $n$, $\alpha\in S_{0}(n)$ and $h_{1},\ldots,h_{n}\in\mathcal{H}_{\mathbb{R}}$, we will consider symbols $\mathnormal{T}\left(\alpha\otimes(h_{1}\otimes\ldots\otimes h_{n})\right)$ subject to the symmetry relation
@@
Here for $\alpha\in S_{0}(n)$, $\beta\in S_{0}(k)$, the permutation $\alpha\cup\beta\in S_{0}(n+k)$ is obtained by shifting the cycles of $\beta$ by $n$, and merging the cycles of $\alpha$ and $\beta$ containing 0. See Notation 3.2.
Next, let $q$ be a non-zero parameter. For a transposition $\tau$, we define the contraction $C_{\tau}$ as follows. For each $n$ and $\alpha\in S_{0}(n)$, $C_{\tau}(\alpha)$ is obtained by: multiplying $\tau\alpha$; erasing the support of $\tau$ and shifting to obtain a permutation in $S_{0}(n-2)$; and multiplying by a weight depending on $q$ and the number of cycles in $\tau\alpha$. See Definition 3.9. Then as usual, for each $n$, we define the Laplacian on $\mathbb{C}[S_{0}(n)]$ as the sum over transpositions $\mathcal{L}=\sum_{\tau\in S(n)}C_{\tau}$, and the Hermite trace polynomial as $\mathnormal{W}\left(\eta\otimes_{s}F\right)=\mathnormal{T}\left(e^{-\mathcal{L}}(\eta\otimes_{s}F)\right)$. Here, the contraction on the tensor part of the argument is the usual tensor contraction. Hermite trace polynomials satisfy several properties which parallel those of ordinary Hermite polynomials.
We may now define a linear functional $\varphi$ on $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ by requiring it to be unital and zero on each $\mathnormal{W}\left(\eta\otimes_{s}F\right)$. Note that with respect to the spanning set $\left\{\mathnormal{T}\left(\eta\otimes_{s}F\right)\right\}$, the coefficients for multiplication do not depend on $q$, while those of the linear functional $\varphi$ do. On the other hand, if we use $\left\{\mathnormal{W}\left(\eta\otimes_{s}F\right)\right\}$ as a spanning set, the coefficients for multiplication depend on $q$, while those of the linear functional $\varphi$ do not. $\varphi$ is positive if and only if $q$ is zero or of the form $\pm\frac{1}{N}$ for $N\in\mathbb{N}$ [16, 28]. For such $q$, we can define the GNS Hilbert space $\mathcal{F}_{q}(\mathcal{H})$ as the completion of the quotient $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$ of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ with respect to the seminorm induced by $\varphi$. The inner product takes the form
@@
Since $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ is naturally graded, we may interpret $\mathcal{F}_{q}(\mathcal{H})$ as a Fock space. There are several natural choices of creation (and annihilation) operators which we describe. The corresponding field operators live in smaller subalgebras of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$. There is an interesting connection with a different Fock space construction from [6] which we describe in some detail. Incidentally, another article exploring the connection between the characters of the symmetric groups and GUE matrices is [28]. We have not elucidated the connection between their work and ours.
@@
We thus may interpret Hermite trace polynomials as elements $\mathnormal{W}\left(\alpha\otimes_{s}(h_{1}\otimes\ldots\otimes h_{n})\right)$ in the algebra $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$ with the linear functional $\varphi$; or the corresponding trace polynomials in formal variables; or (for $q=1/N$) the corresponding trace polynomials in Gaussian random matrices $\left\{X(h_{i})\right\}$. In this last interpretation, we prove that the matrix entries of Hermite trace polynomials are themselves (multivariate) Hermite polynomials in the matrix entries of $\left\{X(h_{i})\right\}$.
To the best of our knowledge, the Hermite trace polynomials $\mathnormal{W}\left(\eta\otimes_{s}F\right)$ are new even in the univariate case $\mathcal{H}=\mathbb{C}$. However, if we further restrict to $\eta\in\mathbb{C}[S(n)]$ (rather than $S_{0}(n)$), the corresponding objects have appeared in the literature. Indeed, denoting $\chi^{\lambda}$ the character of the irreducible representation indexed by the partition $\lambda$, $\mathnormal{W}\left(\chi^{\lambda}\right)$ is closely related to the corresponding Hermite polynomial of matrix argument. In particular, all of these elements are orthogonal with respect to $\varphi$. Moreover, one can form a more general set of characters $\chi^{\lambda,\lambda^{\prime}}$ indexed by pairs of partitions which differ by one box, such that $\left\{\mathnormal{W}\left(\chi^{\lambda,\lambda^{\prime}}\right)\right\}$ form an orthogonal basis for $\mathcal{F}_{1/N}(\mathbb{C})$. This collection of trace polynomials is clearly deserving of additional study.
@@ 2.1. Permutations and partitions
Denote $n]=\left\{1,\ldots,n\right\}$ and $0,n]=\left\{0,1,\ldots,n\right\}$.
@@
Denote by $S_{0}(n)$ the permutations of $[0,n]$. $S(n)$ is a subgroup of $S_{0}(n)$, and so acts on $S_{0}(n)$ by conjugation. The equivalence classes under this action are subsets of the standard conjugacy classes where the number of elements of the cycle containing 0 is preserved. So they are in a natural bijection with number partitions of $n+1$ with a marked element. Equivalently, they are indexed by pairs of partitions $(\lambda,\lambda^{\prime})$, where $\lambda\in\operatorname{\mathrm{Par}}(n)$, $\lambda^{\prime}\in\operatorname{\mathrm{Par}}(n+1)$, and the corresponding diagrams differ by one box.
For $\alpha\in S_{0}(n)$, denote $\operatorname{\mathrm{cyc}}_{0}(\alpha)=\operatorname{\mathrm{cyc}}(\alpha)-1$, where $\operatorname{\mathrm{cyc}}(\alpha)$ is the number of cycles of $\alpha$. In other words, $\operatorname{\mathrm{cyc}}_{0}(\alpha)$ is the number of cycles of $\alpha$ not containing 0. Denote
@@
For $n<m$, we may sometimes identify the element $\alpha\in S_{0}(n)$ with the corresponding element of $S_{0}(m)$ under the natural inclusion $0,n]\subset[0,m]$ (so that the image of $\alpha$ fixes the elements of $[n+1,m]$). Note that $\left|\alpha\right|$ is preserved by this identification.
A partial permutation $\alpha\in\mathcal{PS}_{0}(n)$ is a bijection from a subset of $[0,n]$ onto a subset of $[0,n]$; proper subsets and the empty subset are allowed. Orbits of a partial permutation fall into two types. Cyclic orbits are cycles in the usual permutation sense. Linear orbits have an initial and a final element. Note that a linear orbit has at least two elements. It is convenient to abuse the terminology and consider elements of $[0,n]$ which do not belong to any orbit of $\alpha$ as single-element linear orbits of $\alpha$. Denote $\mathcal{PS}_{0}(n,N)$ the set of partial permutations of $[0,n]$ with $N$ linear orbits.
@@ 2.2. Structure theory of the symmetric group
See [15] for basic representation theory background. For a partition $\lambda\in\operatorname{\mathrm{Par}}(n+1)$, denote by $\chi^{\lambda}$ the character of the irreducible representation of $S_{0}(n)$ corresponding to $\lambda$. We will identify $\chi^{\lambda}$ with the element
@@
[…]
The following are well-known [38], [15].
@@
  • For $\lambda\in\operatorname{\mathrm{Par}}(n)$, write $\lambda^{\prime}=\lambda+\square$ if the Young diagram for $\lambda$ is obtained by removing one of the boxes from the Young diagram for $\lambda^{\prime}$. Denote $\chi^{\lambda^{\prime}:\lambda}$ the character of the compression of the $\lambda^{\prime}$-irreducible representation of $S_{0}(n)$ to the (unique) component giving a $\lambda$-irreducible representation of $S(n)$. Then $\left\{\chi^{\lambda^{\prime}:\lambda}:\lambda\in\operatorname{\mathrm{Par}}(n),\lambda^{\prime}=\lambda+\square\right\}$ are orthogonal and span $Z(\mathbb{C}[S_{0}(n)]:\mathbb{C}[S(n)])$.
@@
See for example [7] for a detailed exposition and explicit expressions for $\mathcal{W}(E_{ij}^{\lambda})$.
@@ 2.4. Gaussian Hilbert spaces
See [20] for proofs and more details.
@@
and $\mathcal{P}(\mathcal{H}_{\mathbb{R}})=\bigcup_{n=0}^{\infty}\mathcal{P}_{n}(\mathcal{H}_{\mathbb{R}})=$ all such polynomials. Denote $\overline{\mathcal{P}}_{n}(\mathcal{H}_{\mathbb{R}})$ the closure of $\mathcal{P}_{n}(\mathcal{H}_{\mathbb{R}})$ in $L^{2}(\mathcal{P}(\mathcal{H}_{\mathbb{R}}),\mathbb{E})$ and $\mathcal{H}_{n}=\overline{\mathcal{P}}_{n}(\mathcal{H}_{\mathbb{R}})\ominus\overline{\mathcal{P}}_{n-1}(\mathcal{H}_{\mathbb{R}})$, the orthogonal complement taken in $L^{2}(\mathcal{P}(\mathcal{H}_{\mathbb{R}}),\mathbb{E})$. Denote $p\mapsto:\!\!p\!\!:$ the orthogonal projection from each $\overline{\mathcal{P}}_{n}(\mathcal{H}_{\mathbb{R}})$ onto $\mathcal{H}_{n}$; note that it maps each $\mathcal{P}_{n}(\mathcal{H}_{\mathbb{R}})$ to itself. Define the Wick product map $W$ from the algebraic symmetric Fock space $\bigoplus_{n=0}^{\infty}\mathcal{H}_{\mathbb{R}}^{\odot_{s}n}$ into $L^{2}(\mathcal{P}(\mathcal{H}_{\mathbb{R}}),\mathbb{E})$ by the linear extension of
@@
(2) $\mathnormal{W}\left(f_{1}\otimes\ldots\otimes f_{n}\right)=\sum_{\pi\in\mathcal{P}_{1,2}(n)}(-1)^{\left|\pi\right|}\prod_{\left\{i,j\right\}\in\mathrm{Pair}\left(\pi\right)}\left\langle f_{i},f_{j}\right\rangle\prod_{k\in\mathrm{Sing}\left(\pi\right)}X(f_{k}).$
@@
In the complex case, there are two natural types of Gaussian Hilbert spaces.
  • We may take $\mathcal{H}=\left\{f+ig:f,g\in\mathcal{H}_{\mathbb{R}}\right\}$ to be the complexification of $\mathcal{H}_{\mathbb{R}}$, with $X(f+ig)=X(f)+iX(g)$. Then each $X(h)$ is centered complex Gaussian, and their covariance is $\mathbb{E}[X(h)X(k)]=\left\langle h,\bar{k}\right\rangle_{\mathcal{H}}$.
@@
If $\left\{\xi_{i}:i\in\Xi\right\}$ is an orthonormal basis for $\mathcal{H}_{\mathbb{R}}$, then
$\left\{\mathnormal{W}\left(\xi_{u(1)},\ldots,\xi_{u(n)}\right):n\geq 0,\mathbf{u}\in\Xi^{n}\right\}$
is an orthogonal basis for $L^{2}(\mathcal{P}(\mathcal{H}_{\mathbb{R}}),\mathbb{E})$. Similarly,
$\left\{Z(\xi_{u(1)})\ldots Z(\xi_{u(n)}):n\geq 0,\mathbf{u}\in\Xi^{n}\right\}$
is an orthogonal basis for $L^{2}(\mathcal{P}(\mathcal{K}),\mathbb{E})$. Thus the linear extension of the map $\mathcal{S}:\mathnormal{W}\left(h_{1},\ldots,h_{n}\right)\mapsto Z(h_{1})\ldots Z(h_{n})$ (the Segal-Bargmann transform) is an isomorphism between these spaces.